OpenStudy (anonymous):

(sqrt 7x^7)(sqrt 14x^9y^9

OpenStudy (anonymous):

\[\sqrt{7xy ^{7}}\sqrt{14x ^{9}y ^{9}}\]

OpenStudy (jack1):

(sqrt 7xy^7)(sqrt 14x^9y^9) so sqrt = to the power of 1/2 so write like that to begin with

OpenStudy (jack1):

(sqrt 7xy^7)(sqrt 14x^9y^9) = (7^1/2 * x^1/2 * y^7/2 ) times (14^1/2) * x^9/2 * y^9/2) now add like terms using the law of powers, example: x^y * x^z = x^(y+z)

OpenStudy (jack1):

(sqrt 7xy^7)(sqrt 14x^9y^9) = (7^1/2 * x^1/2 * y^7/2 ) times (14^1/2) * x^9/2 * y^9/2) now (14^1/2) = 2 (7^1/2) so (7^1/2 * x^1/2 * y^7/2 ) times (14^1/2) * x^9/2 * y^9/2) = (7^1/2 * 2 * 7^1/2 * x^1/2 * x^9/2 * y^7/2 * y^9/2) = 2 (7^1/2 * 7^1/2 * x^1/2 * x^9/2 * y^7/2 * y^9/2) = 2 [ 7^(1/2 + 1/2) * x^ (1/2 + 9/2) * y^(7/2 + 9/2) ] are you ok to simplify from here @cronus8992 ?

OpenStudy (anonymous):

I keep getting \[7x ^{10}y ^{8}\sqrt{2}\] Which I know is wrong

OpenStudy (jack1):

ok, 1'm getting sqrt2 * 7^1 * x^10/2 * y^16/2 = sqrt2 * 7 * x^5 * y^8

OpenStudy (anonymous):

I combined them like this \[\sqrt{(7xy ^{10})(14x^9y^9)}\] With that I get \[\sqrt{98x ^{20}y ^{16}}\] From there I am stuck

OpenStudy (jack1):

shouldnt be x^20

OpenStudy (jack1):

so 98 = 49 * 2 sqrt 98 = sqrt 49 * sqrt 2 = 7 * sqrt 2 if that helps

OpenStudy (anonymous):

why would x^10 be wrong? x(x^19) right

OpenStudy (anonymous):

x^20 sorry

OpenStudy (jack1):

youve got x^1 * x^9 = x^(1+9) = x^10

OpenStudy (anonymous):

Oh simple mistype thank you so much that actually fixed the problem lol idk why i input 20

OpenStudy (jack1):

cools u all good then?